On January 1 of the current year, the Queen Corporation issued 12% bonds with a face value of $57,000. The bonds are sold for $55,290. The bonds pay interest semiannually on June 30 and December 31 and the maturity date is December 31, five years from now. Queen records straight-line amortization of the bond discount. Determine the bond interest expense for the year ended December 31.

Select the correct answer.

$7,182

$570

$6,840

$1,710
Answer #1
Semi-annual cash interest 3420 =57000*12%*6/12
Semi-annual discount amortization 171 =(57000-55290)/5*6/12
Semi-annual interest expense 3591
Bond interest expense for the year ended December 31 7182 =3591+3591
Option A $7,182 is correct